Reports 1-1 of 1 Clear search Modify search
takafumi.ushiba - 15:39 Sunday 01 May 2022 (20632) Print this report
In-air health check of ETMY: check of the spectra of sensors

Similar work with klog20631.

I checked the spectra of sensors on ETMY.
BF LVDT signals have larger noise at high frequencies than ITMY's, which seem similar problem that we found in ITMX GAS before (klog18789).
I also found that MNH1 and IMV3 spectra has several times larger noise than ITMY's.
In addition, there are small jumps of signals (~ 1cnt) of MNH2 though it is not appears in spectra.
They should be investigated before vacuum closing.

Figure 1 to 5 show the spectra of sensors on IP, BF, GAS, MN, IM stages, respectively (same graphs posted in klog20631).
Followings are what I noticed from the spectra.
1. all BF LVDTs of ETMY have larger noise than ITMY's at high frequencies.
2. MNH1 photosensor has larger noise than the others
3. IMV3 photosensors also have larger noise.

In addition to this, I found the strange jump of MNH2 photosensor signals in time series data (fig6).
The jump is small (~ 1cnt) and we cannot see any excess noise in spectrum.
Since the jump is very small, it is unlikely to be a problem of sensr itself but ADC problem as we faced before like klog20197 and klog19212.
Anyway, we need to investigate it.

We used spectra of REFNUM 0025 for this check.
Data were stored at /kagra/Dropbox/Measurements/VIS/SPECTRA/0/

Images attached to this report
Comments to this report:
ryutaro.takahashi - 17:59 Monday 09 May 2022 (20666) Print this report

[Takahashi, Sato]

We replaced the main card of the LVDT driver for the BF damper to reduce the noise level. The main card is embedded in the center of the driver chassis (photo). The old card (S1807754) was replaced with the new card (S2113543). The noise floor level was improved from 0.02~0.09 to 0.01 at 10Hz.

Images attached to this comment
takafumi.ushiba - 21:00 Tuesday 10 May 2022 (20684) Print this report

[Yamamoto, Ushiba]

ETMY photosensors' spectra were checked again after ADC noise investigation.
Then, we found many strange peak stractures and large excess of several photosensors, which are obtained in previous spectra measurement, are coming from ADC.
Though MNH1 and IMV3 photosensors seem to have slightly larger noise, they look acceptable.

After ADC noise investigation (klog20682), we checked photosensor noise by using relatively better noise performance channels of ADC (ADC1 channel 3-7).
Figure 1 shows the result.
Black line is the ADC noise level of each channel, and other lines are photosensor noise (see legend for confirming which photosensor noises are shown).

Figure 2 shows the photosensor noises at MN stage (top) and IM stage (bottom).
Followings are some remarkable notes:
1. MNH1 photosensor has slightly larger noise than the other MN photosensors between 2Hz and 40Hz.
2. IMV3 photosensor has larger noise than the other IM photosensors between 0.8Hz and 100 Hz.
3. Below 0.1Hz, almost all photsensor noises are limited by AC noise.

For 1, it doesn't seem a big problem because the excess noise is very small (less than a factor of 2).
For 2, it also doesn't a big problem because IM photosensor won't be used for local control.
Tomorrow, I will continue to investigate where these noise differences are coming but no large hardware trouble was not found.

Measured data was stored at /users/VISsvn/TypeApayload/ETMY/Spectra/20220510/.

Images attached to this comment
Search Help